# Smoothing-Enabled Randomized Stochastic Gradient Schemes for Solving Nonconvex Nonsmooth Potential Games under Uncertainty

### v1

@@ 1 Introduction
The Nash equilibrium (NE) [45] serves as a central solution concept in noncooperative multi-agent decision-making, capturing situations in which no participant can reduce its cost or improve its payoff by unilaterally altering its strategy. This framework naturally arises in engineered fields and economic environments [13, 22, 42, 57, 58] where rational agents interact through coupled constraints or shared resources. The study of NE has therefore become essential for characterizing system-level outcomes that emerge from decentralized behavior. In recent years, the increasing focus on large-scale optimization and machine learning has further expanded the role of Nash equilibria, such as convex–concave saddle-point minimax problems [34, 39, 62] in high-dimensional regimes. Many efforts in computing equilibria in continuous-strategy games with both smooth and nonsmooth objectives have relied on gradient-response and best-response methods, along with their stochastic variants, largely under the setting where each player’s problem is convex [16, 32, 36, 37, 38]. In this paper, we consider the setting where player-specific functions are possibly nonconvex, nonsmooth, and expectation-valued. Specifically, our focus lies on the stochastic $N$-player noncooperative game
@@ 2 Preliminaries
Given $x\in X\subseteq\text{dom}\,f$ and $d\in\mathbb{R}^{n}$, we denote the generalized directional derivative of $f$ at $x$ in the direction $d$ by $f^{\circ}(x;d)$. Let $\partial^{C}f(x)$ denote the Clarke subdifferential of $f$ at $x$ if $f$ is Lipschitz continuous. We denote the Fréchet normal cone to $X$ at $x$ by $\mathcal{N}_{X}(x)$. We refer the reader to the monographs [7, 11, 55] for their definitions and properties.
@@ 2.2 Clarke-Nash Equilibria and Generalized Variational Inequalities
Inspired by Clarke stationarity [7, Definition 6.1.4], Clarke-Nash equilibrium (CNE) were introduced. We recall the definition of CNE below.
Definition 2.1 ([7, Definition 11.1.1]).
@@
which can be captured by the solution set of $\mathrm{VI}\>(X,F)$, where $X\triangleq\prod_{i=1}^{N}X_{i}$ and $F(x)\triangleq(\nabla_{x_{i}}f_{i}(x))_{i=1}^{N}$. The equilibria existence under smoothness from the VI perspective is established in [63, Theorem 1].
@@
By Definition 2.1, we know that the solution set of $\mathrm{GVI}\>(X,\mathbf{F})$ characterizes CNE. Indeed, we may take $\mathbf{F}(x)\triangleq(\partial^{C}_{x_{i}}f_{i}(x_{i},x_{-i}))_{i=1}^{N}$. The following proposition establishes the CNE existence from the perspective of $\mathrm{GVI}\>(X,\mathbf{F})$, under the upper semicontinuity [7, Definition 1.5.1] assumption on $\mathbf{F}$.
@@
The proof proceeds similarly to [7, Proposition 11.2.2] by considering the set-valued map $\mathbf{F}(x)\triangleq(\partial^{C}_{x_{i}}f_{i}(x_{i},x_{-i}))_{i=1}^{N}$. The required upper semicontinuity in [7, Proposition 11.2.2] of Clarke subdifferentials follows from [7, Proposition 4.3.1] directly. ∎
@@ 3 Stochastic Nonconvex Smooth Potential Games
The following proposition shows that under smoothness, a potential game can be equivalently viewed as an optimization problem. In fact, our claim is equivalent to [43, Lemma 4.4] and [56, Definition 3.1.1], which are often used as an alternative definition of potential games.
@@ 3.1 Randomized Stochastic Gradient Scheme
One interesting result is that, when the potential function $P$ is smooth and pseudoconvex, RSG converges to a NE, even though the game is nonconvex. We formalize it in the following theorem and its proof follows from [63, Proposition 3] immediately.
@@ 4.1 Randomized Smoothed RSG Scheme
where the penultimate inequality is due to the unbiasedness $\mathbb{E}[\Delta^{k,h}_{i,l}\,|\,x^{k}_{i}]=\nabla_{x_{i}}h^{\eta}_{i}(x^{k}_{i})$ and the last inequality is from [41, Lemma 2.4-(v)]. ∎
@@ 4.2 Approximation of CNE
We note that neither convexity nor closedness is required in Lemma 4.2. Based on it, we next show that under the Lipschitz continuity of $\partial^{C}h_{i}(x_{i})$, the one-sided deviation $\mathbb{D}(\partial^{C}_{\eta}h_{i}(x_{i}),\partial^{C}h_{i}(x_{i}))$ can be upper bounded, where $\partial^{C}_{\eta}h_{i}(x_{i})$ is the $\eta$-Clarke subdifferential defined in (1). Recall that a set-valued mapping $F:X\rightrightarrows Y$ is said to be $L$-Lipschitz on $X$ [55, Definition 9.26] if there exists $L>0$ such that $F(x_{1})\subseteq F(x_{2})\,+\,L\|x_{1}-x_{2}\|\mathbb{B}_{Y}$ for any $x_{1},x_{2}\in X$, where $\mathbb{B}_{Y}$ is the unit ball in $Y$.
@@
for any $i\in[N]$. For any $\zeta_{i}\in\partial^{C}_{\eta}h_{i}(x_{i})$, it follows from the convex combination expression of convex hull (see [44, Proposition 1.30]) that
@@
Recall from [7, Proposition 4.3.1] that the Clarke subdifferential is a convex compact set. Therefore, the projection of $\nabla_{x_{i}}h^{\eta}_{i}(x^{R}_{i})$ onto $\partial^{C}h_{i}(x^{R}_{i})$ is unique. We denote it by
@@
where the penultimate inequality is implied by invoking [19, Proposition 1]. By noting the projection equality (29), we may bound the term $\|u^{R}-(\nabla_{x_{i}}h^{\eta}_{i}(x^{R}_{i}))_{i=1}^{N}\|^{2}$ in (4.2) by
@@ 5.1 Biased RS-RSG Scheme
The proof can be found in [10, Theorem 2]. We omit the proof here. ∎
@@
By [41, Lemma 2.4-(v)] and the fact that $\mathbb{E}[\Delta^{k,h}_{i,l}\,|\,x^{k}]=\nabla_{x_{i}}\mathbb{E}[\tilde{h}^{\eta}_{i}(x^{k}_{i},y_{i}(x^{k}_{i}),\xi)]$, we obtain
@@ 7 Conclusions
where the last inequality follows from [19, Proposition 1]. Summing up the above inequalities for $k=1,\dots,T$ and noting that $0<\gamma_{k}\leq 1/L<2/L$ with $0<\gamma_{k}<1/L$ for at least one $k$, we obtain
@@
Then, by [19, Proposition 1], we have from the above inequality and (56) that

### v2

@@ 1 Introduction
The Nash equilibrium (NE) [43] serves as a central solution concept in noncooperative multi-agent decision-making, capturing situations in which no participant can reduce its cost or improve its payoff by unilaterally altering its strategy. This framework naturally arises in engineered fields and economic environments [20, 40, 55, 56] where rational agents interact through coupled constraints or shared resources. The study of NE has therefore become essential for characterizing system-level outcomes that emerge from decentralized behavior. In recent years, the increasing focus on large-scale optimization and machine learning has further expanded the role of Nash equilibria, such as convex–concave saddle-point minimax problems [32, 37, 60] in high-dimensional regimes. Many efforts in computing equilibria in continuous-strategy games with both smooth and nonsmooth objectives have relied on gradient-response and best-response methods, along with their stochastic variants, largely under the setting where each player’s problem is convex [14, 30, 35, 36, 34]. In this paper, we consider the setting where player-specific functions are possibly nonconvex, nonsmooth, and expectation-valued. Specifically, our focus lies on the stochastic $N$-player noncooperative game
@@ 2 Preliminaries
Given $x\in X\subseteq\text{dom}\,f$ and $d\in\mathbb{R}^{n}$, we denote the generalized directional derivative of $f$ at $x$ in the direction $d$ by $f^{\circ}(x;d)$. Let $\partial^{C}f(x)$ denote the Clarke subdifferential of $f$ at $x$ if $f$ is Lipschitz continuous. We denote the Fréchet normal cone to $X$ at $x$ by $\mathcal{N}_{X}(x)$. We refer the reader to the monographs [10, 53] for their definitions and properties.
@@ 2.2 Clarke-Nash Equilibria and Generalized Variational Inequalities
Inspired by Clarke stationarity [10, Definition 6.1.4], Clarke-Nash equilibrium (CNE) were introduced. We recall the definition of CNE below.
Definition 2.1 ([10, Definition 11.1.1]).
@@
which can be captured by the solution set of $\mathrm{VI}\>(X,F)$, where $X\triangleq\prod_{i=1}^{N}X_{i}$ and $F(x)\triangleq(\nabla_{x_{i}}f_{i}(x))_{i=1}^{N}$. The equilibria existence under smoothness from the VI perspective is established in [61, Theorem 1].
@@
By Definition 2.1, we know that the solution set of $\mathrm{GVI}\>(X,\mathbf{F})$ characterizes CNE. Indeed, we may take $\mathbf{F}(x)\triangleq(\partial^{C}_{x_{i}}f_{i}(x_{i},x_{-i}))_{i=1}^{N}$. The following proposition establishes the CNE existence from the perspective of $\mathrm{GVI}\>(X,\mathbf{F})$, under the upper semicontinuity [10, Definition 1.5.1] assumption on $\mathbf{F}$.
@@
The proof proceeds similarly to [10, Proposition 11.2.2] by considering the set-valued map $\mathbf{F}(x)\triangleq(\partial^{C}_{x_{i}}f_{i}(x_{i},x_{-i}))_{i=1}^{N}$. The required upper semicontinuity in [10, Proposition 11.2.2] of Clarke subdifferentials follows from [10, Proposition 4.3.1] directly. ∎
@@ 3 Stochastic Nonconvex Smooth Potential Games
The following proposition shows that under smoothness, a potential game can be equivalently viewed as an optimization problem. In fact, our claim is equivalent to [41, Lemma 4.4] and [54, Definition 3.1.1], which are often used as an alternative definition of potential games.
@@ 3.1 Randomized Stochastic Gradient Scheme
One interesting result is that, when the potential function $P$ is smooth and pseudoconvex, RSG converges to a NE, even though the game is nonconvex. We formalize it in the following theorem and its proof follows from [61, Proposition 3] immediately.
@@ 4.1 Randomized Smoothed RSG Scheme
where the penultimate inequality is due to the unbiasedness $\mathbb{E}[\Delta^{k,h}_{i,l}\,|\,x^{k}_{i}]=\nabla_{x_{i}}h^{\eta}_{i}(x^{k}_{i})$ and the last inequality is from [39, Lemma 2.4-(v)]. ∎
@@ 4.2 Approximation of CNE
We note that neither convexity nor closedness is required in Lemma 4.2. Based on it, we next show that under the Lipschitz continuity of $\partial^{C}h_{i}(x_{i})$, the one-sided deviation $\mathbb{D}(\partial^{C}_{\eta}h_{i}(x_{i}),\partial^{C}h_{i}(x_{i}))$ can be upper bounded, where $\partial^{C}_{\eta}h_{i}(x_{i})$ is the $\eta$-Clarke subdifferential defined in (1). Recall that a set-valued mapping $F:X\rightrightarrows Y$ is said to be $L$-Lipschitz on $X$ [53, Definition 9.26] if there exists $L>0$ such that $F(x_{1})\subseteq F(x_{2})\,+\,L\|x_{1}-x_{2}\|\mathbb{B}_{Y}$ for any $x_{1},x_{2}\in X$, where $\mathbb{B}_{Y}$ is the unit ball in $Y$.
@@
for any $i\in[N]$. For any $\zeta_{i}\in\partial^{C}_{\eta}h_{i}(x_{i})$, it follows from the convex combination expression of convex hull (see [42, Proposition 1.30]) that
@@
Recall from [10, Proposition 4.3.1] that the Clarke subdifferential is a convex compact set. Therefore, the projection of $\nabla_{x_{i}}h^{\eta}_{i}(x^{R}_{i})$ onto $\partial^{C}h_{i}(x^{R}_{i})$ is unique. We denote it by
@@
where the penultimate inequality is implied by invoking [16, Proposition 1]. By noting the projection equality (29), we may bound the term $\|u^{R}-(\nabla_{x_{i}}h^{\eta}_{i}(x^{R}_{i}))_{i=1}^{N}\|^{2}$ in (4.2) by
@@ 5.1 Biased RS-RSG Scheme
The proof can be found in [8, Theorem 2]. We omit the proof here. ∎
@@
By [39, Lemma 2.4-(v)] and the fact that $\mathbb{E}[\Delta^{k,h}_{i,l}\,|\,x^{k}]=\nabla_{x_{i}}\mathbb{E}[\tilde{h}^{\eta}_{i}(x^{k}_{i},y_{i}(x^{k}_{i}),\xi)]$, we obtain
@@ 7 Conclusions
where the last inequality follows from [16, Proposition 1]. Summing up the above inequalities for $k=1,\dots,T$ and noting that $0<\gamma_{k}\leq 1/L<2/L$ with $0<\gamma_{k}<1/L$ for at least one $k$, we obtain
@@
Then, by [16, Proposition 1], we have from the above inequality and (56) that